# What Does Stable Diffusion Know about the 3D Scene?

### v2

@@ 3.3 Property Probing
We aim to determine which Stable Diffusion features best represent different properties. To obtain features from an off-the-shelf Stable Diffusion network, we follow the approach of Tang et al. (2023) used for DIFT, where noise is added to the input image in the latent space, and features are extracted from different layers and time steps of the model. While probing the properties, linear classifiers are used to infer the relationships between regions. The region representation is obtained by a simple average pooling of the diffusion features over the annotated region or object.
@@
where $f_{\theta_{l}}$ is the $l$-th U-Net layer. In this way, we can get the representation of an image $F_{t,l}$ at time step $t$ and $l$-th U-Net layer for the probe. We upsample the obtained representation to the size of original image with bi-linear, then use the region mask to get a region-wise feature vector, by averaging the feature vectors of each pixel it contains, i.e., average pooling.
@@
where $v_{k,t,l}$ is the feature vector of the $k$-th region $R_{k}$. $\odot$ here is a per-pixel product of the region mask and the feature.
Linear Probing. After computing features from a diffusion model, we use a linear classifier (a linear SVM) to examine how well these features can be used to answer questions to each of the properties. Specifically, the input of the classifier is the difference or absolute difference between the feature vectors of Region A and Region B, i.e., $v_{A}-v_{B}$ or $|v_{A}-v_{B}|$, and the output is a Yes/No answer to the question. Denoting the answer to the question as $Q$, then since the questions about Same Plane, Perpendicular Plane, Material, Shadow and Occlusion are symmetric relations, $Q(v_{A},v_{B})=Q(v_{B},v_{A})$. However, the questions about Support Relation and Depth are not symmetric. Thus, we use $|v_{A}-v_{B}|$ (a symmetric function) as input for the first group of questions, and $v_{A}-v_{B}$ (non-symmetric) for the rest of questions. We train the linear classifier on the train set via the positive/negative samples of region pairs for each property; do a grid search on the validation set to find (i) the optimal time step $t$, (ii) the U-Net layer $l$, and (iii) the SVM regularization parameter $C$; and evaluate the performance on the test set.
Discussion. Some of the current symmetric questions can be reformulated in a non-symmetric manner in order to obtain more information about the property. For example, the shadow question could be formulated as ‘is region A the shadow of object B’ rather than the (symmetric) ‘are region A and region B in an object-shadow relationship’. The non-symmetric version requires the classifier to explicitly identify which region is the object, and which the shadow. Note, the protocol has been explained for diffusion models, but in the following we also apply it to other networks trained at large-scale. In Section 4.3 we give results for its application to OpenCLIP (Radford et al., 2021; Ilharco et al., 2021), DINOv1 (Caron et al., 2021), DINOv2 (Oquab et al., 2023), and VQGAN (Esser et al., 2021).
@@ 4 Experiments
4.1 Implementation Detail and Evaluation Metric
Implementation Details. For each property, we sample the same number of positive / negative pairs, to maintain a balanced evaluation set. In terms of the linear SVM, we tune the penalty parameter $C$ on the val split to find the best $C$ for each property. Therefore, we are grid searching 3 parameters on the val set, namely, Stable Diffusion Timestep $t$ ranging from 1 to 1000, U-Net Layer $l$ covering the 4 downsampling and 4 upsampling layers, and the SVM penalty parameter $C$ ranging over $0.001,0.01,0.1,1,10,100,1000$. The timestep is searched with a stride of 20 steps, since the difference in performance around the optimal value varies by less than 0.01 within 20 steps. In practice the $C$ parameter is always between $0.1$ and $1$, so we carry out a finer search over values between $0.1$ and $1.0$ in steps of $0.1$. The linear SVM is solved using the libsvm library (Chang & Lin, 2011) with the SMO algorithm, to get the unique global optimal solution. Please refer to the appendix for more implementation details.
@@ 4.1 Implementation Detail and Evaluation Metric
Table 3: SVM grid search results of other features trained at scale. We train the linear SVM on the training set, and grid search the best combination of ViT/Transformer layer and $C$ on the validation set. The OpenCLIP and VQGAN models we use have 48 transformer layers, DINOv1 has 12 layers and DINOv2 has 40 layers. The $i$-th layer means the $i$-th transformer layer from the input side.
@@
Table 4: Performance of Stable Diffusion features compared to state-of-the-art self-supervised features. For each property, we use the best time step, layer and $C$ found in the grid search in Table 2 for Stable Diffusion, and the best layer and $C$ found in the grid search for other features. The performance is the ROC AUC on the test set, and ‘Random’ means a random classifier.
@@
4.2 Results for Stable Diffusion
The results for grid search are shown in Table 2. For Stable Diffusion U-Net Layer, $D_{l}$ means the $l$-th layer of the U-Net decoder, i.e., upsampling layer, from outside to inside (right to left), and we provide an illustration of the layers in Figure 5(a). We can make two observations: First, the best time step for different properties varies, while the best U-Net layer is always $D_{3}$ in the decoder rather than the encoder. Further explorations using Stable Diffusion features for downstream tasks could thus start from the U-Net decoder layers, especially $D_{3}$. Second, in terms of the performance on the test set, we find that Stable Diffusion can understand very well about scene geometry, support relations, shadows, and depth, but it is less performant at predicting material and occlusion. Examples of its failure are shown in Figure 5 (b) (c). As noted in Zhan et al. (2022) and Kirillov et al. (2023), grouping all separated parts of an object due to occlusion remains challenging even for state-of-the-art detection and segmentation models.
4.3 Results for Other Features Trained at Scale
We have also applied our protocol to other models pre-trained on large scale image datasets, including OpenCLIP (Radford et al., 2021; Ilharco et al., 2021) pre-trained on LAION dataset (Schuhmann et al., 2022), DINOv1 (Caron et al., 2021) pre-trained on ImageNet dataset (Deng et al., 2009), DINOv2 (Oquab et al., 2023) pre-trained on LVD-142M dataset (Oquab et al., 2023), and VQGAN (Esser et al., 2021) pre-trained on ImageNet dataset (Deng et al., 2009). We use the best pre-trained checkpoints available on their official GitHub – ViT-B for DINOv1, ViT-G for OpenCLIP and DINOv2, and the 48-layer transformer checkpoint for VQGAN. Similar to Stable Diffusion, for each of these models, we conduct a grid search on the validation set in terms of the ViT/Transformer layer and $C$ for SVM, and use the best combination of parameters for evaluation on the test set. Grid search results are reported in Table 3 for the tasks of material and support. Results of other tasks are in the appendix.
4.4 Comparison of Different Features Trained at Scale
We compare the state-of-the-art pre-trained large-scale models’ representations on various downstream tasks in Table 4. It can be observed that the Stable Diffusion and DINOv2 representations outperform OpenCLIP, DINOv1 and VQGAN for all properties, indicating the potential of utilizing Stable Diffusion and DINOv2 representations for different downstream tasks with the optimal time steps and layers found in Section 4.2 and Section 4.3.
@@ 5 Discussion and Future Work
In this paper, we have developed a protocol to examine whether the Stable Diffusion and other models pre-trained on large scale image datasets, like CLIP, DINO and VQGAN, have explicit feature representations for different properties of the 3D scene.
It is interesting to find that different time steps of Stable Diffusion and different layers of DINOv2 representations can handle several different physical properties at a state-of-the-art performance, indicating the potential of utilising the Stable Diffusion and DINOv2 models for different downstream tasks. However, for some properties such as material and occlusion, the networks are not distilling the information in a manner that can be used by a linear probe. This could indicate that these properties are not modelled well by the network or that more than a linear probe is required to tease them out. We show examples of the prediction failures for these properties in Figure 5. In the appendix, we show that such prediction failures also occur in generated (i.e. synthetic) images. It is worth noting that occlusion is a challenge even for the powerful Segment Anything Model (SAM) (Kirillov et al., 2023), where the model ‘hallucinates small disconnected components at times’.
This paper has given some insight into answering the question: ‘What does Stable Diffusion know about the 3D scene?’ for real images. Of course, there are more properties that could be investigated in the manner proposed here. For example, contact relations (Fouhey et al., 2016) and object orientation (Xiang et al., 2018), as well as the more nuanced non-symmetric formulations of the current questions. Another direction would be to use the pixel-wise supervision method of Bhattad et al. (2023) to search for features that can predict depth maps, normal maps, etc. Furthermore, utilising combinations of features from different time steps and layers as in Luo et al. (2023); Zhang et al. (2023) could be explored to obtain even better performance for using Stable Diffusion features for downstream tasks.
@@ Appendix A More Implementation Details
Following DIFT (Tang et al., 2023), when we extract Stable Diffusion features, we add a different random noise 8 times and then take the average of the generated features. The process can be completed in one forward pass as we are using a batch of 8. We use an empty prompt ‘’ as the text prompt.
@@
For the partition of train/val split, we select the train & val images from different scenes for the NYUv2 (Silberman et al., 2012) and ScanNetv2 (Dai et al., 2017) dataset.
@@
As there are some noisy annotations in the (Liu et al., 2019) dataset, we manually filter the images whose annotations are inaccurate.
@@
To study the occlusion problem, we utilise the Separated COCO dataset (Zhan et al., 2022). The original dataset only collects separated objects due to occlusion in the COCO 2017 val split, we further extend it to the COCO 2017 train split for more data using the same method as in (Zhan et al., 2022). As the original COCO val split is too small (only 5k images v.s. 118k images in the COCO train split), we get our partition of train/val/test splits by dividing the generated dataset ourselves.
Appendix B Analysis of Stable Diffusion Generated Images
As Figure 1 shows, our motivation for the paper is that we observe that Stable Diffusion correctly predicts different physical properties of the scene. The reason why we do not study the generated images directly is that there are no annotations available on different properties for these synthetic images, so it is expensive to get quantitative results. But in this section, we provide more qualitative examples and analysis of Stable Diffusion generated images in terms of different physical properties. The observations match our findings on real images – Stable Diffusion ‘knows’ about a number of physical properties including scene geometry, material, support relations, shadows, occlusion and depth, but may fail in some cases in terms of material and occlusion.
We show examples for: Scene Geometry in Figure 6; Material, Support Relations, and Shadows in Figure 7; and Occlusion and Depth in Figure 8.
@@ Appendix B Analysis of Stable Diffusion Generated Images
Appendix C Additional Results for Other Features Trained at Large Scale
As mentioned in Section 4.3, we have conducted grid search for all the other large pre-trained models, including OpenCLIP, DINOv1, DINOv2 and VQGAN for all tasks. Tables in this section provide results for the Same Plane (Table 5), Perpendicular Plane (Table 6), Shadow (Table 7), Occlusion (Table 8) and Depth (Table 9) tasks for these models.

### v3

@@ 3.3 Property Probing
Take Stable Diffusion as an example, we aim to determine which features best represent different properties. To obtain features from an off-the-shelf Stable Diffusion network, we follow the approach of [38] used for DIFT, where noise is added to the input image in the latent space, and features are extracted from different layers and time steps of the model. While probing the properties, linear classifiers are used to infer the relationships between regions. The region representation is obtained by a simple average pooling of the diffusion features over the annotated region or object.
@@
where $f_{\theta_{l}}$ is the $l$-th U-Net layer. In this way, we can get the representation of an image $F_{t,l}$ at time step $t$ and $l$-th U-Net layer for the probe. We upsample the obtained representation to the size of the original image with bi-linear, then use the region mask to get a region-wise feature vector, by averaging the feature vectors of each pixel it contains, i.e., average pooling.
@@
where $v_{k,t,l}$ is the feature vector of the $k$-th region $R_{k}$. $\odot$ here is a per-pixel product of the region mask and the feature. For other models, including CLIP, DINOv1, DINOv2 and VQGAN, we feed the image into the ViT/Transformer and extract features from different layers. Then use average pooling as in Equation 3 to obtain the feature for each region.
Linear Probing. After extracting features from large-scale vision models, we use a linear classifier (a linear SVM) to examine how well these features can be used to answer questions to each of the properties. Specifically, the input of the classifier is the difference or absolute difference between the feature vectors of Region A and Region B, i.e., $v_{A}-v_{B}$ or $|v_{A}-v_{B}|$, and the output is a Yes/No answer to the question. Denoting the answer to the question as $Q$, then since the questions about Same Plane, Perpendicular Plane, Material, Shadow and Occlusion are symmetric relations, $Q(v_{A},v_{B})=Q(v_{B},v_{A})$. However, the questions about Support Relation and Depth are not symmetric. Thus, we use $|v_{A}-v_{B}|$ (a symmetric function) as input for the first group of questions, and $v_{A}-v_{B}$ (non-symmetric) for the rest of questions. We train the linear classifier on the train set via the positive/negative samples of region pairs for each property; do a grid search on the validation set to find (i) the optimal time step $t$ (only for Stable Diffusion), (ii) the U-Net layer $l$ for Stable Diffusion and the Transformer layer $l$ for other models, and (iii) the SVM regularization parameter $C$; and evaluate the performance on the test set. The grid search is only feasible because our proposed protocol is lightweight, and can assess the effectiveness of features for different downstream tasks with minimal resource demands.
@@ 4 Experiments
4.1 Implementation Details and Evaluation Metric
Implementation Details. For each property, we sample the same number of positive / negative pairs, to maintain a balanced evaluation set. In terms of the linear SVM, we tune the penalty parameter $C$ on the val split to find the best $C$ for each property. Therefore, we are grid searching 3 parameters on the val set, namely, Timestep $t$ ranging from 1 to 1000 (only for Stable Diffusion), U-Net Layer $l$ covering the 4 downsampling and 4 upsampling layers for Stable Diffusion and Transformer Layer $l$ for other networks, and the SVM penalty parameter $C$ ranging over $0.001,0.01,0.1,1,10,100,1000$. The timestep is searched with a stride of 20 steps, since the difference in performance around the optimal value varies by less than 0.01 within 20 steps. In practice the $C$ parameter is always between $0.1$ and $1$, so we carry out a finer search over values between $0.1$ and $1.0$ in steps of $0.1$. The linear SVM is solved using the libsvm library [6] with the SMO algorithm, to get the unique global optimal solution. Please refer to the appendix for more implementation details.
@@ 4.2 Results for Stable Diffusion
The results for grid search are shown in Table 2. For Stable Diffusion U-Net Layer, $D_{l}$ means the $l$-th layer of the U-Net decoder, i.e., upsampling layer, from outside to inside (right to left), and we provide an illustration of the layers in Figure 5(a). We can make observations: First, the best time step for different properties varies, but the optimal time step is usually before 400. This is expected as a large time step adds too much noise, so not much information is contained about the image. Second, in terms of the layer, the best U-Net layer is always $D_{3}$ in the decoder rather than the encoder. The optimal layer is in the middle, as D1 is too close to the noise space and D4 has just started decoding. Further explorations using Stable Diffusion features for downstream tasks could thus start from the U-Net decoder layers, especially $D_{3}$. Third, in terms of the performance on the test set, we find that Stable Diffusion can understand very well about scene geometry, support relations, shadows, and depth, but it is less performant at predicting material and occlusion. Examples of its failure are shown in Figure 5 (b) (c). As noted in [48] and [22], grouping all separated parts of an object due to occlusion remains challenging even for state-of-the-art detection and segmentation models. The appendix gives grid search results at all time steps and layers.
4.3 Results for CLIP/DINO/VQGAN Features
Table 3: SVM grid search results of CLIP/DINO/VQGAN features. We train the linear SVM on the training set, and grid search the best combination of ViT/Transformer layer and $C$ on the validation set. The OpenCLIP and VQGAN models we use have 48 transformer layers, DINOv1 has 12 layers and DINOv2 has 40 layers. The $i$-th layer means the $i$-th transformer layer from the input side.
@@ 4.3 Results for CLIP/DINO/VQGAN Features
In this section we show grid search results for OpenCLIP [30, 18] pre-trained on LAION dataset [33], DINOv1 [5] pre-trained on ImageNet dataset [9], DINOv2 [27] pre-trained on LVD-142M dataset [27], and VQGAN [11] pre-trained on ImageNet dataset [9]. We use the best pre-trained checkpoints available on their official GitHub – ViT-B for DINOv1, ViT-G for OpenCLIP and DINOv2, and the 48-layer transformer checkpoint for VQGAN. Similar to Stable Diffusion, for each of these models, we conduct a grid search on the validation set in terms of the ViT/Transformer layer and $C$ for SVM, and use the best combination of parameters for evaluation on the test set.
Grid search results are reported in Table 3 for the tasks of material and support. It can be observed that different layers of different models are good at different properties. Results of other tasks are in the appendix.
4.4 Comparison of Different Features Trained at Scale
Table 4: Comparison of different features trained at scale. For each property, we use the best time step, layer and $C$ found in the grid search for Stable Diffusion, and the best layer and $C$ found in the grid search for other features. The performance is the ROC AUC on the test set, and ‘Random’ means a random classifier.
@@ 4.4 Comparison of Different Features Trained at Scale
We compare the state-of-the-art pre-trained large-scale vison models’ representations on various downstream tasks in Table 4. It can be observed that the Stable Diffusion and DINOv2 representations outperform OpenCLIP, DINOv1 and VQGAN for all properties, indicating the potential of utilizing Stable Diffusion and DINOv2 representations for different downstream tasks with the optimal time steps and layers found in Section 4.2 and Section 4.3.
@@ 5 Discussion and Future Work
In this paper, we have developed a general and lightweight protocol to efficiently examine whether models pre-trained on large scale image datasets, like CLIP, DINO, VQGAN and Stable Diffusion, have explicit feature representations for different properties of the 3D physical scene.
It is interesting to find that different time steps of Stable Diffusion and different layers of DINOv2 representations can handle several different physical properties at a state-of-the-art performance, indicating the potential of utilising the Stable Diffusion and DINOv2 models for different downstream tasks.
However, for some properties such as material and occlusion, the networks are not distilling the information in a manner that can be used by a linear probe. This could indicate that these properties are not modelled well by the network or that more than a linear probe is required to tease them out. We show examples of the prediction failures for these properties in Figure 5. In the appendix, we show that such prediction failures also occur in generated (i.e. synthetic) images. It is worth noting that occlusion is a challenge even for the powerful Segment Anything Model (SAM) [22], where the model ‘hallucinates small disconnected components at times’.
In the appendix, we provide preliminary results of using the probed Stable Diffusion feature for downstream tasks (Section F). We also provide examples of another use case of spotting Stable Diffusion generated images based on the properties that the model is not good at (Section B).
This paper has given some insight into answering the question: ‘To what extent do large vision models understand the 3D scene’ for real images. Of course, there are more properties that could be investigated in the manner proposed here. For example, contact relations [14] and object orientation [45], as well as the more nuanced non-symmetric formulations of the current questions.
@@ Appendix A More Implementation Details
Following DIFT [38], when we extract Stable Diffusion features, we add a different random noise 8 times and then take the average of the generated features. The process can be completed in one forward pass as we are using a batch of 8. We use an empty prompt ‘’ as the text prompt.
@@
For the partition of train/val split, we select the train & val images from different scenes for the NYUv2 [36] and ScanNetv2 [8] dataset.
@@
As there are some noisy annotations in the [24] dataset, we manually filter the images whose annotations are inaccurate.
@@
To study the occlusion problem, we utilise the Separated COCO dataset [48]. The original dataset only collects separated objects due to occlusion in the COCO 2017 val split, we further extend it to the COCO 2017 train split for more data using the same method as in [48]. As the original COCO val split is too small (only 5k images v.s. 118k images in the COCO train split), we get our partition of train/val/test splits by dividing the generated dataset ourselves.
Computing Resources.
We run experiments on CPU with 20G memory for SVM training/testing.
Appendix B More Stable Diffusion Generated Images
Here we give more examples of Stable Diffusion generated images as mentioned in the caption of Figure 1. We show examples for: Scene Geometry in Figure 6; Material, Support Relations, and Shadows in Figure 7; and Occlusion and Depth in Figure 8.
The observations match our findings on studying the Stable Diffusion features – Stable Diffusion ‘knows’ about a number of physical properties including scene geometry, material, support relations, shadows, occlusion and depth, but may fail in some cases in terms of material and occlusion.
As mentioned in Section 5, we can spot Stable Diffusion generated images via the properties it is not good at. Take the second row of Figure 7 and the first row of Figure 8 as examples – we can spot that the images are generated by Stable Diffusion because of the failure to generate a clear boundary between two different materials, and the failure to connect separated parts due to occlusion.
@@ Appendix B More Stable Diffusion Generated Images
Appendix C Additional Results for CLIP/DINO/VQGAN Features Trained at Large Scale
As mentioned in Section 4.3, we have conducted grid search for OpenCLIP, DINOv1, DINOv2 and VQGAN for all tasks. Tables in this section provide results for the Same Plane (Table 5), Perpendicular Plane (Table 6), Shadow (Table 7), Occlusion (Table 8) and Depth (Table 9) tasks for these models.
@@ Appendix D Train/Val AUC Results of Stable Diffusion Features
Appendix E AUC Curves for Stable Diffusion Features Grid Search
As mentioned in Section 4.2, we provide curves for AUC at different layers and time steps of probing Stale Diffusion features in Figure 9 - Figure 15.
Figure 9: Curves for AUC at different layers and time steps of probing Stable Diffusion for the same plane task.
Figure 10: Curves for AUC at different layers and time steps of probing Stable Diffusion for the perpendicular plane task.
Figure 11: Curves for AUC at different layers and time steps of probing Stable Diffusion for the material task.
Figure 12: Curves for AUC at different layers and time steps of probing Stable Diffusion for the support task.
Figure 13: Curves for AUC at different layers and time steps of probing Stable Diffusion for the shadow task.
Figure 14: Curves for AUC at different layers and time steps of probing Stable Diffusion for the occlusion task.
Figure 15: Curves for AUC at different layers and time steps of probing Stable Diffusion for the depth task.
Appendix F Preliminary Results of Applying Probed Feature for Downstream Tasks
As mentioned in Section 5, we provide preliminary results of using our probed feature for downstream task in this section. Here we take the task of surface normal estimation as an example. We have conducted experiments on the applications of using the certain time step and layer of Stable Diffusion features as we probed (here using the probed feature of the ‘perpendicular plane’ task). In particular, we directly injected our extracted features into the open-source state-of-the-art model iDisc [29] at where the encoder encodes the image feature. Without bells and whistles, when injecting Stable Diffusion feature and trained for only 1000 iterations (they train 45000 iterations in total), the model performs better on most of the metrics compared with their reported number (as in Table 11), which indicates our proposed feature probing strategy is significant for the corresponding 3D tasks. Additionally, beyond traditional 3D tasks, our study provides a wider image of 3D physical understanding and can also inspire methods to improve other physical tasks, including Instance Shadow Detection [41], Material Segmentation [39], Support Relation Inference [36], if our probed features are injected or utilised, which has not been well-studied before.
Table 11: Preliminary results of using the probed feature for downstream task. Here we show the results of injecting our probed Stable Diffusion feature to iDisc [29]. Please see text for more details.
Model Mean Angular Error $\downarrow$ Angular RMSE $\downarrow$ % < 11.25 $\uparrow$ % < 22.5 $\uparrow$ % < 30 $\uparrow$
iDisc 14.6 22.8 63.8 79.8 85.6
Ours 13.8 18.8 58.1 80.9 88.7